OpenStudy (goformit100):

What universal set(s) would you propose for each of the following : (i) The set of right triangles. (ii) The set of isosceles triangles

OpenStudy (goformit100):

@LovingMyself

OpenStudy (goformit100):

HELP^

OpenStudy (anonymous):

Set of triangles...is d universal set...for both to be subsets.

OpenStudy (anonymous):

Set of right triangles : Is a set such that its one angle is 90 degrees.. or if its sides are a,b,c then a^2= b^2 + c^2 (a= hypotenuse , Pyrhagoras theorem)

OpenStudy (anonymous):

Isosceles Triangle : Is a set such that if its sides are a,b,c then ab=ac and angle b= angle c ..

OpenStudy (anonymous):

Hope that Helps :)

OpenStudy (goformit100):

Thanks

OpenStudy (anonymous):

Np